Three positive charges are located in the x-y planCan anyone help me with the second part
Answer #1

Refer to the diagram

01 02 3:6 cm -1 .-....-箕灯ir.ei仲 . Fx B 3cn -5 4 3 2 1 0 1 2 3 4 5

In triangle ABC

AB = 2 cm, BC = 3 cm ,

AC= \sqrt{2^{2}+3^{2}}= 3.6 cm

Angle BAC \angle BAC = \tan^{-1}\left ( \frac{3}{2} \right )= 56.3^{0}

Now Force on Q3  due to Q2

F = \frac{kQ_{2}Q_{3}}{(AC)^{2}} = \frac{9\times 10^{9}\times 5.20\times 10^{-6}\times 6.30\times 10^{-6}}{(3.6\times 10^{-2})^{2}}= 227.5 N

Now the x component of the F is given by

F_{x} =F\sin 56.3^{0}= 227.5 N\times \sin 56.3^{0} = 189.27 N